How to solve this equation please:
0.5(c+2.8) - c = 0.6c+0.3
Step by step would be nice. Anything is appreciated.

Answer by checkley77(12844)

0.5(c+2.8)-c=0.6c+0.3
.5c+1.4-c=.6c+.3
.5c-c-.6c=.3-1.4
-1.1c=-1.1
c=-1.1/-1.1
c=1 ans.
Proof:
.5(1+2.8)-1=.6*1+.3
.5*3.8-1=.6+.3
1.9-1=-.9
.9=.9